OpenStudy (anonymous):

At equilibrium, what is the hydrogen-ion concentration if the acid dissociation constant is 0.000 001 and the acid concentration is 0.01M?

OpenStudy (joannablackwelder):

\[K _{a}=\frac{ \left[ products \right] }{ \left[ reactants \right] }\]

OpenStudy (joannablackwelder):

So, to know what the products and reactants are, we will need to write the equilibrium reaction.

OpenStudy (joannablackwelder):

Making sense so far, @fashionlove?

OpenStudy (anonymous):

yes, what is the nest step?

OpenStudy (joannablackwelder):

Write a general acid dissociation reaction: HA <-> H+ + A-

OpenStudy (joannablackwelder):

Can you write the Ka equation?

OpenStudy (joannablackwelder):

Sweet! :)

OpenStudy (joannablackwelder):

Plug in your numbers and solve for x.

OpenStudy (joannablackwelder):

Over 0.01, but I think that is what you meant

OpenStudy (joannablackwelder):

0.000001*0.01=x^2

OpenStudy (joannablackwelder):

That means 1x10^-8

OpenStudy (joannablackwelder):

Right!

OpenStudy (joannablackwelder):

Which simplifies to..?

OpenStudy (joannablackwelder):

Yup! Good job!

OpenStudy (anonymous):

thanks:)

OpenStudy (joannablackwelder):

No worries :)